OpenStudy (natsent06):

5% of a number is 35. What is that number?

OpenStudy (natsent06):

who are you talking to

OpenStudy (anonymous):

0.5 * x = 35 x = 35 / 0.5 x = 70

OpenStudy (anonymous):

0.5 is 50% 0.05*x=35 x=35/0.05 x=700

OpenStudy (natsent06):

-_- really i just answered that question and i cant go back to it

OpenStudy (anonymous):

@breezymeetee all your fault. smh.

OpenStudy (anonymous):

I answered it right type in a calculator 0.5 * 70 = 35 I got it right! @logancoop

OpenStudy (anonymous):

But 0.5 is 50% not 5% :p